OpenStudy (anonymous):

Find an equation of the tangent line to the curve y=8^x at the point (2,64) Tangent line y=

OpenStudy (anonymous):

Find the first derivative to find the slope at that point. y'=ln(8) * 8^x

OpenStudy (anonymous):

once you have the slope you can find the rest of the equation using y=mx+b. Use your point for y and x and solve or b

OpenStudy (anonymous):

So would I plug in 64 for y and 2 for x into the first derivative to get the slope? Which would make it 2.0794....5?

OpenStudy (anonymous):

the first derivative can also be written like this \[dy/dt=\ln(8)8^x\] use the x value to find the slope.

OpenStudy (anonymous):

dy/dx sorry

OpenStudy (anonymous):

the slope would be 133.08426

OpenStudy (anonymous):

alright so I got 133.084=ln(8)*8^2 I now go 64=133.084(2)+b to find b?

OpenStudy (anonymous):

you got it!

OpenStudy (anonymous):

Thank You so much. I appreciate it